Hello, see problem below
The following data represent the number of children in a sample of 10 families from Chicago: 4, 2, 1, 1, 5, 3, 0, 1, 0, and 2.
Compute the variance of the data. Place your answer, rounded to two decimal places, in the blank.
I have tried this problem more than one time and it getting 2.49 and i was told it should be 2.77.
These are the steps I took to complete it:
The sum of the values is 19 and the total number of values is 10 that gives me a mean of 1.9. Then I subtracted the mean from each value then took the square root of the result. Then I added all the squared results and I got 24.9, divided that by 10 to get the variance of 2.49
If I missed something, please let know.

Answer by rothauserc(4718)

variance is the average of the squared differences from the mean
:
first calculate the mean
:
mean = (4 +2 +1 +1 +5 +3 +0 +1 +0 +2)/10 = 1.9
:
variance = (4-1.9)^2 +(2-1.9)^2 +(1-1.9)^2 +(1-1.9)^2 +(5-1.9)^2 +(3-1.9)^2 +(0-1.9)^2 +(1-1.9)^2 +(0-1.9)^2 +(2-1.9)^2 / 10 = 2.49
:
you are correct
: